# 3D sans 3D Scans: Scalable Pre-training from Video-Generated Point Clouds

### v1

@@ 5 Experiments
VGGT [wang2025vggt] 49.3 75.1
MapAnything [keetha2025mapanything] 50.1 75.5
@@ 5.1 Implementation Details
Pre-training configuration. We pre-train the model using LAM3C with a Point Transformerv3 (PTv3) [Wu_2024_CVPR] backbone. Each scene provides 9D input features consisting of coordinates, colors and normals. Training is conducted on eight NVIDIA H200 GPUs with a total batch size of 16. AdamW [loshchilov2017decoupled] is used as the optimizer with an initial learning rate of 0.001 and layer-wise decay of 0.9. The OneCycleLR [smith2018disciplined] learning rate scheduler is used, running for 145,600 iterations with a warm-up period followed by cosine annealing. Weight decay is linearly increased from 0.04 to 0.1 during pre-training.
Evaluation protocol. We evaluate the pre-training effectiveness of LAM3C on semantic and instance segmentation tasks in indoor scenes. We use four widely used indoor scene datasets for evaluation: ScanNet [dai2017scannet], ScanNet++ [yeshwanth2023scannet++], ScanNet200 [rozenberszki2022language], and S3DIS [Armeni_2016_CVPR]. Following standard evaluation protocols, we report results under two settings: (i) Full fine-tuning, in which all pre-trained model parameters are updated on the fine-tuning dataset, and (ii) Linear probing, in which the backbone parameters are frozen and only a final linear layer is trained on the fine-tuning dataset. The details of the hyperparameters for each dataset are described in the supplementary material.
@@ 5.2 Main results
Semantic segmentation. Here, we evaluate the effect of pre-training on downstream semantic segmentation tasks using mIoU as the evaluation metric. The results are shown in Table 1. First, compared to training PTv3 from scratch, LAM3C improves performance across all datasets under both linear probing and full fine-tuning settings. This shows that effective representations can be acquired even through pre-training using only video-generated point clouds. In addition, LAM3C trained with video-generated point clouds and real point clouds achieved equivalent or better performance across all datasets compared to the baseline models Sonata (all real) and Sonata (ScanNet).
This suggests that the pseudo data generated by LAM3C provides complementary information rather than simply replacing real data. Based on these results, LAM3C can be considered an effective pre-training method for semantic segmentation tasks in indoor scenes.
Instance segmentation. This paper also evaluates the pre-training effect on instance segmentation tasks. AP is used as the evaluation metric, and the results are shown in Table 2. First, compared to training PTv3 from scratch, LAM3C consistently improves AP across all datasets under both linear probing and full fine-tuning settings. This demonstrates that representations enabling individual object recognition can be acquired even through pre-training without real 3D scans. Next, we show that LAM3C trained with real point clouds achieves higher performance than video generated point cloud only and comparable mAP compared to the baseline models Sonata (all real) and Sonata (ScanNet).
This suggests that the video-generated point clouds provides complementary information regarding the geometric distinctiveness required for instance segmentation, rather than merely substituting real data. Therefore, LAM3C can be considered an effective pre-training method for instance segmentation in indoor scenes.
@@ 5.3 Ablation studies
Effect of LAM3C components (see Table LABEL:tab:self-distillation and Table LABEL:tab:regularizer_loss). This experiment verifies the effectiveness of noise-regularized loss of LAM3C. First, we compare the results of using only the clustering loss with those of adding our proposed noise regularized loss. Table LABEL:tab:self-distillation confirms that noise regularized loss achieves an improvement in segmentation performance compared to clustering loss alone, for both LP and Full-FT. This suggests that noise-regularized loss is effective for RoomTours composed of video-generated point clouds containing noise and missing regions, and helps stabilize representation learning.
Furthermore, we investigate the combined effects of the Laplacian smoothing loss and the noise consistency loss in LAM3C. Table LABEL:tab:regularizer_loss shows that noise-regularized loss using both loss functions achieves higher pre-training effectiveness than using either Laplacian smoothing loss or noise consistency loss alone. This result indicates that the Laplacian smoothing loss and the noise consistency loss play complementary roles in promoting stable learning on noisy point clouds.
Effect of data scaling (see Table LABEL:tab:data_scaling). RoomTours can generate video-generated point clouds from videos without human work, making it easier to scale than real 3D scans, provided that video and computational resources are available. This experiment investigates the effect of pre-training on the number of video-generated point cloud scenes in RoomTours. Compared to 1k scenes, increasing the number of scenes to 16k and 50k tends to improve performance. These results demonstrate that LAM3C, despite not using real 3D scans, holds potential for scaling with video-generated point clouds.
Effect of reconstruction model (see Table LABEL:tab:reconstruct-model). We assess the impact of feed-forward reconstruction models on video-generated point clouds. This experiment compares VGGT, MapAnything, and $\pi^{3}$. Table LABEL:tab:reconstruct-model shows that $\pi^{3}$ exhibits greater pre-training effectiveness than VGGT and MapAnything. This is because $\pi^{3}$ can process longer sequences, enabling it to reconstruct high-quality video-generated point clouds compared with VGGT and MapAnything.
@@ 5.4 Additional Experiments
Impact of limited resources. In this experiment, we evaluate the effect of LAM3C under limited resorces for downstream tasks. On ScanNet semantic segmentation, we restrict the number of training scenes to {1%, 5%, 10%, 40%}, and separately limit the number of annotated point clouds to {20, 50, 100, 200}. As shown in Table 4, our method provides clear gains even when both the number of training samples and the amount of supervision are heavily constrained. For all settings except for 1% and 5%, LAM3C achieves higher segmentation accuracy than Sonata trained with real 3D scans. The drop in very small-data regimes can be attributed to the domain gap between real point clouds and video-generated point clouds.
@@
[…]
* uses PTv3 (Large) and 434k pre-training steps.
Qualitative results. This experiment qualitatively evaluates zero-shot recognition by using PCA. As shown in Fig. 6, LAM3C produces clear segmentation of local structures such as desks even in a zero-shot setting. This result indicates that video-generated point clouds preserve enough coarse geometric cues to learn representations for 3D-SSL directly from video. At the same time, LAM3C exhibits less coherent global structure compared to Sonata trained with real 3D scans. For example, floor edges show lower contrast, and the floor boundaries appear more blurred in the PCA space. This behavior is expected because $\pi^{3}$ reconstructions provide approximate 3D representations with variations in coordinate frames and scale, which makes it more difficult for the model to learn globally consistent scene geometry than when the model is pre-trained on accurate 3D scans.
6 Discussion
Our findings show that high-fidelity 3D scans are not a prerequisite for effective 3D-SSL. Despite being noisy, incomplete, and inconsistent, video-generated point clouds reconstructed from unlabeled videos prove sufficient for 3D-SSL. Perhaps most notably, LAM3C performs strongly in linear probing, a setting that has not been widely explored in previous 3D-SSL methods. This indicates that these imperfect reconstructions still encode rich geometic cues relevant to real-world tasks. Overall, our results highlight a promising direction toward scalable 3D data generation. Instead of costly real 3D scans, unlabeled videos offer a far more scalable source of 3D supervision, where the key is learning to handle noisy and missing regions effectively.
Figure 6: Zero-shot qualitative evaluation using PCA visualizations. This experiments show that LAM3C learns semantic representations in real world without pre-training with real 3D scans.
@@ 7 Conclusion
We have presented LAM3C, a self-supervised framework that learns representations from unlabeled videos. By leveraging video-generated point clouds reconstructed via a feed-forward reconstruction model and incorporating noise-regularized loss comprising local smoothness and noise consistency, LAM3C achieves equal or better pre-training performance without any real 3D scans. We constructed RoomTours, a dataset of 49k video-generated point clouds derived from unlabeled room-tour videos, and demonstrated that LAM3C matches or surpasses prior pre-training methods such as Sonata even under limited downstream data conditions. These results suggest that unlabeled videos can serve as an effective data source for scalable 3D-SSL.
Limitations and outlook. Although the reconstruction models used in this paper use real point clouds during training, our core contribution lies in showing that previously untapped resources, such as web videos can be made useful for 3D-SSL and improve with scale. By piggybacking on reconstruction models, our approach benefits from the fast progress in this domain and might enable unifying 3D reconstruction and recognition models.
@@ Acknowledgment
We used ABCI 3.0 provided by AIST with support from “ABCI 3.0 Development Acceleration Use”.
Figure 7: Statistics of the RoomTours video collection. Our RoomTours consists of 3,462 indoor walkthrough videos collected from 19 countries. (Left) Geographic distribution of the source countries and their proportions. (Right) Distribution of video durations, ranging from short clips to long walkthroughs, with a median length of 3.63 minutes.
Figure 8: Statistics of the video-generated point clouds after our alignment procedure. We report the properties of the aligned scenes used for pretraining. (Left) Distribution of the number of points per scene. (Center) Distribution of scene scales (the diagonal length of the axis-aligned bounding box). (Right) Distribution of the median point spacing, defined as the median kNN distance within each scene.
@@ Appendix A RoomTours Details
We construct RoomTours using a feedforward reconstruction model applied to unlabeled videos from the web. The dataset consists of video-generated point clouds obtained from both our collected videos and existing datasets (RealEstate10k [zhou2018stereo], YouTube House Tours [chang2020semantic], and HouseTours [Celen_2025_ICCV]). In this section, we report statistics of our independently collected videos and the video-generated point clouds reconstructed from them.
For our own collection, we assume that indoor layouts and furniture vary across geographic regions. Therefore, we collected walkthrough videos from 19 countries, as shown in Fig. 7. In total, we gathered 3,462 videos, each with an median duration of 3.63 minutes.
For scene classification, we sample every frame at the native FPS of each video and resize all frames to the standard CLIP [radford2021learning] ViT-B/32 resolution of $224\times 224$. Each frame is first classified by the CLIP image encoder as either “indoor” or “outdoor.” Frames predicted as indoor are then re-evaluated using room-type prompts (living room, bedroom, bathroom) to obtain scene categorization. This two-stage filtering produces a total of 15,921 indoor sequences, with each video contributing on average 4.59 sequences.
The reconstruction of a single scene takes approximately 5 minutes on average, although the actual time depends on the number of frames extracted from each video. We process our collected videos using eight NVIDIA H200 GPUs (144GB each), requiring roughly eight hours in total to generate the video-based point clouds. The raw outputs are pixel-aligned point clouds and therefore contain an extremely large number of points. Directly applying our post-processing pipeline to these raw point clouds leads to substantial computational overhead. To maintain efficiency, we randomly downsample each scene to 20k points before alignment and normalization. Fig. 8 reports the statistics of the post-processed video-generated point clouds, including the distribution of point counts and scene scale. After post-processing, the structural properties of the indoor scenes closely match those observed in standard ScanNet scenes.
@@
iterations 291,200 (RoomTours-50k, PTv3-Base)
436,800 (RoomTours-50k, PTv3-Large)
@@ Appendix B Implementation Details
This section explains the details of our implementation. We begin by describing Point Transformer V3 (PTv3) [Wu_2024_CVPR] which is the backbone model used our experiments. We then detail the pre-training setup and the configurations adopted for semantic segmentation and instance segmentation. This paper used PointCept [pointcept2023] for the implementation as base code.
Backbone network. PTv3 improves efficiency by removing the kNN query and the Relative Positional Encoding (RPE) occupy modules in PTv2 [wu2022point]. The kNN query module was designed to capture local structure on unstructured point clouds, but it requires repeated distance computations at every iteration and occupies 28% of the forward time. The RPE occupy module encodes relative positions, yet point cloud RPE must compute pairwise Euclidean distances for every point pair, accounting for 26% of the computation. PTv3 avoids both costs by abandoning the unordered-set formulation and serializing point clouds into a structured representation, which eliminates the need for kNN and RPE occupy and leads to efficiency gains. PTv3 assigns a consistent order to points through four ordering strategies: Z-order (Z), Transposed Z-order (TZ), Hilbert (H), and Transposed Hilbert (TH).
We use PTv3 as the backbone model for all experiments, with two variants: PTv3 (Base) and PTv3 (Large). The Base variant follows the configuration used in Sonata, and the Large variant follows the configuration used in Concerto. The detailed parameter settings are provided in Table 5.
@@
Table 9: Hyperparamter Analysis Experiments on the ScanNet semantic segmentation. We report Full Fine-tuning (Full-FT) and Linear Probing (LP) performance (%). Unless noted otherwise, the default configuration uses RoomTours-1k generated by $\pi^{3}$ and is pre-trained with PTv3 (Base). Both Full-FT and LP are trained for 100 epochs. Default settings are highlighted in gray.
pre-training setting. We set the default parameters as shown in Table 6. The Laplacian normalization loss hyperparameters were set based on experimental results and computational considerations during pre-training. Other settings related to the clustering loss follow the configuration used in Sonata. Furthermore, our comparative experiments employed three settings: RoomTours-16k (PTv3-Base), RoomTours-50k (PTv3-Base), and RoomTours-50k (PTv3-Large). Based on empirical results from previous research, the number of iterations is scaled with both data size and model size
Downstream task setting. We evaluate the pre-training effect on two downstream tasks: semantic segmentation and instance segmentation in indoor scenes. Semantic segmentation and instance segmentation are evaluated on ScanNet [dai2017scannet], ScanNet200 [rozenberszki2022language], ScanNet++ [yeshwanth2023scannet++], S3DIS [Armeni_2016_CVPR]. For both tasks, we adopt the following fine-tuning protocols: Linear probing and Full fine-tuning. Linear probing freezes the entire backbone and trains only a linear classifier, whereas Full fine-tuning updates all model parameters. For the linear probing and the full fine-tuning protocol, we follow Sonata setting [wu2025sonata]. The configuration for semantic segmentation is provided in Table 7, and the configuration for instance segmentation is provided in Table 8.

### v2

@@ 5 Experiments
VGGT [45] 49.3 75.1
MapAnything [21] 50.1 75.5
@@ 5.1 Implementation Details
Pre-training configuration. We pre-train the model using LAM3C with a Point Transformer V3 (PTv3) [50] backbone. Each scene provides 9D input features consisting of coordinates, colors and normals. Training is conducted on eight NVIDIA H200 GPUs with a total batch size of 16. AdamW [25] is used as the optimizer with an initial learning rate of 0.001 and layer-wise decay of 0.9. The OneCycleLR [40] learning rate scheduler is used, running for 145,600 iterations with a warm-up period followed by cosine annealing. Weight decay is linearly increased from 0.04 to 0.1 during pre-training. For the default RoomTours-1k / PTv3-Base setting, we train for 145,600 iterations with OneCycleLR. For larger-scale settings, we scale the number of iterations with dataset and model size (see Supp. Tab. 7).
Evaluation protocol. We evaluate the pre-training effectiveness of LAM3C on semantic and instance segmentation tasks in indoor scenes. We use four widely used indoor scene datasets for evaluation: ScanNet [15], ScanNet++ [56], ScanNet200 [35], and S3DIS [2]. Following standard evaluation protocols, we report results under two settings: (i) Full fine-tuning, in which all pre-trained model parameters are updated on the fine-tuning dataset, and (ii) Linear probing, in which the backbone parameters are frozen and only a final linear layer is trained on the fine-tuning dataset. The details of the hyperparameters for each dataset are described in the supplementary material.
@@ 5.2 Main results
Semantic segmentation. Here, we evaluate the effect of pre-training on downstream semantic segmentation tasks using mIoU as the evaluation metric. The results are shown in Table 1. First, compared to training PTv3 from scratch, LAM3C improves performance across all datasets under both linear probing and full fine-tuning settings. This shows that effective representations can be acquired even through pre-training using only VGPC.
This suggests that the VGPC provide complementary information rather than simply replacing real data. Based on these results, LAM3C can be considered an effective pre-training method for semantic segmentation tasks in indoor scenes.
Instance segmentation. We also evaluate the pre-training effect on instance segmentation tasks. AP is used as the evaluation metric, and the results are shown in Table 2. First, compared to training PTv3 from scratch, LAM3C consistently improves AP across all datasets under both linear probing and full fine-tuning settings. This demonstrates that representations enabling individual object recognition can be acquired even through pre-training without real 3D scans. This suggests that VGPC provide complementary information regarding the geometric distinctiveness required for instance segmentation, rather than merely substituting real data. Therefore, LAM3C can be considered an effective pre-training method for instance segmentation in indoor scenes.
@@ 5.3 Ablation studies
Effect of LAM3C components (see Table LABEL:tab:self-distillation and Table LABEL:tab:regularizer_loss). This experiment verifies the effectiveness of noise-regularized loss of LAM3C. First, we compare the results of using only the clustering loss with those of adding our proposed noise-regularized loss. Table LABEL:tab:self-distillation confirms that noise-regularized loss achieves an improvement in segmentation performance compared to clustering loss alone, for both LP and Full-FT. This suggests that noise-regularized loss is effective for RoomTours composed of VGPC containing noise and missing regions, and helps stabilize representation learning. Furthermore, we investigate the combined effects of the Laplacian smoothing loss and the noise consistency loss in LAM3C. Table LABEL:tab:regularizer_loss shows that noise-regularized loss using both loss functions achieves higher pre-training effectiveness than using either Laplacian smoothing loss or noise consistency loss alone. This result indicates that the Laplacian smoothing loss and the noise consistency loss play complementary roles in promoting stable learning on noisy point clouds.
Effect of data scaling (see Table LABEL:tab:data_scaling). RoomTours enables scalable VGPC generation from videos with limited manual intervention, making it easier to scale than real 3D scans, provided that video and computational resources are available. This experiment investigates the effect of pre-training on the number of VGPC scenes in RoomTours. Compared to 1k scenes, increasing the number of scenes to 16k and 49k tends to improve performance. These results demonstrate that LAM3C, despite not using real 3D scans, holds potential for scaling with VGPC.
Effect of reconstruction model (see Table LABEL:tab:reconstruct-model). We assess the impact of feed-forward reconstruction models on VGPC. This experiment compares VGGT, MapAnything, and $\pi^{3}$. Table LABEL:tab:reconstruct-model shows that $\pi^{3}$ exhibits greater pre-training effectiveness than VGGT and MapAnything. This is because $\pi^{3}$ can process longer sequences, enabling it to reconstruct high-quality VGPC compared with VGGT and MapAnything.
@@ 5.4 Additional Experiments
Impact of limited resources. In this experiment, we evaluate the effect of LAM3C under limited resources for downstream tasks. On ScanNet semantic segmentation, we restrict the number of training scenes to {1%, 5%, 10%, 20%}, and separately limit the number of annotated point clouds to {20, 50, 100, 200}. As shown in Table 4, our method provides clear gains even when both the number of training samples and the amount of supervision are heavily constrained. LAM3C remains competitive under limited supervision and consistently outperforms the ScanNet only Sonata baseline, while approaching or exceeding stronger Sonata variants in several settings.
@@
* uses PTv3 (Large) and 437k pre-training steps.
Table 5: Synergy between RoomTours and LAM3C
Sonata LAM3C
ScanNet (clean) 67.1 66.8
RoomTours (noisy) 51.1 57.1
Disentangling the contributions of LAM3C and RoomTours. When pre-trained on ScanNet, the two self-supervised methods, Sonata and LAM3C, perform comparably, indicating that our proposed self-supervised method does not provide inherent gains on curated 3D scans, as shown in Table 5. In contrast, on RoomTours-1k, LAM3C substantially outperforms Sonata, indicating that the improvement arises from the learning design rather than the dataset alone. Taken together, RoomTours provides scalable scan-free pre-training data, while LAM3C is essential for learning effectively from noisy reconstructed point clouds.
Qualitative results. This experiment qualitatively evaluates zero-shot recognition by using PCA. As shown in Fig. 6, LAM3C produces clear segmentation of local structures such as desks even in a zero-shot setting. This result indicates that VGPC preserve enough coarse geometric cues to learn representations for 3D-SSL directly from video. At the same time, LAM3C exhibits less coherent global structure compared to Sonata trained with real 3D scans. For example, floor edges show lower contrast, and the floor boundaries appear more blurred in the PCA space. This behavior is expected because $\pi^{3}$ reconstructions provide approximate 3D representations with variations in coordinate frames and scale, which makes it more difficult for the model to learn globally consistent scene geometry than when the model is pre-trained on accurate 3D scans.
Figure 6: Zero-shot qualitative evaluation using PCA visualizations. LAM3C learns semantic representations in the real world without pre-training on real 3D scans.
@@ 6 Conclusion
We have presented LAM3C, a self-supervised framework that learns 3D representations from unlabeled videos. By leveraging VGPC reconstructed with a feed-forward reconstruction model and introducing noise-regularized objectives, namely local smoothness and noise consistency, LAM3C achieves comparable or better pre-training performance without using any real 3D scans. We also constructed RoomTours, a dataset of 49k VGPC derived from unlabeled room-tour videos, and showed that LAM3C matches or surpasses prior pre-training methods such as Sonata, even under limited downstream data conditions. These results demonstrate that unlabeled videos can serve as an effective and scalable source of supervision for 3D-SSL. Our contribution is this scan-free and scalable 3D-SSL setting based on video-only data collection, rather than the input format itself. Unlike prior 3D-SSL methods, we generate pre-training data from unlabeled videos at scale.
@@ Acknowledgment
This work was supported by the AIST policy-based budget project “R&D on Generative AI Foundation Models for the Physical Domain.” It was also supported by the JSPS Overseas Research Fellowship. In addition, this work was supported by the Japan Science and Technology Agency (JST) through the Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE) program under Grant Number JPMJAP2518. We used ABCI 3.0, provided by AIST and AIST Solutions, with support from the “ABCI 3.0 Development Acceleration Use” program.
@@ Appendix A RoomTours Details
We construct RoomTours using a feedforward reconstruction model applied to unlabeled videos from the web. The dataset consists of VGPC obtained from both our collected videos and existing datasets (RealEstate10k [62], YouTube House Tours [10], and HouseTours [9]). In this section, we report statistics of our independently collected videos and the VGPC reconstructed from them.
For our own collection, we assume that indoor layouts and furniture vary across geographic regions. Therefore, we collected walkthrough videos from 19 countries, as shown in Fig. 7. In total, we gathered 3,462 videos, each with a median duration of 3.63 minutes.
For scene classification, we sample every frame at the native FPS of each video and resize all frames to the standard CLIP [34] ViT-B/32 resolution of $224\times 224$. Each frame is first classified by the CLIP image encoder as either “indoor” or “outdoor.” Frames predicted as indoor are then re-evaluated using room-type prompts (living room, bedroom, bathroom) to obtain scene categorization. This two-stage filtering produces a total of 15,921 indoor sequences, with each video contributing on average 4.59 sequences.
The reconstruction of a single scene takes approximately 5 minutes on average, although the actual time depends on the number of frames extracted from each video. We process our collected videos using eight NVIDIA H200 GPUs (144GB each), requiring roughly eight hours in total to generate the video-based point clouds. The raw outputs are pixel-aligned point clouds and therefore contain an extremely large number of points. Directly applying our post-processing pipeline to these raw point clouds leads to substantial computational overhead. To maintain efficiency, we randomly downsample each scene to 20k points before alignment and normalization. Fig. 8 reports the statistics of the post-processed VGPC, including the distribution of point counts and scene scale. After post-processing, the structural properties of the indoor scenes closely match those observed in standard ScanNet scenes.
@@
iterations 291,200 (RoomTours-49k, PTv3-Base)
436,800 (RoomTours-49k, PTv3-Large)
@@ Appendix B Implementation Details
This section explains the details of our implementation. We begin by describing Point Transformer V3 (PTv3) [50] which is the backbone model used in our experiments. We then detail the pre-training setup and the configurations adopted for semantic segmentation and instance segmentation. Our implementation is based on PointCept [13].
Backbone network. PTv3 improves efficiency by removing the kNN query and the Relative Positional Encoding (RPE) modules in PTv2 [51]. The kNN query module was designed to capture local structure on unstructured point clouds, but it requires repeated distance computations at every iteration and occupies 28% of the forward time. The RPE module encodes relative positions, yet point cloud RPE must compute pairwise Euclidean distances for every point pair, accounting for 26% of the computation. PTv3 avoids both costs by abandoning the unordered-set formulation and serializing point clouds into a structured representation, which eliminates the need for kNN and RPE computations and leads to efficiency gains. PTv3 assigns a consistent order to points through four ordering strategies: Z-order (Z), Transposed Z-order (TZ), Hilbert (H), and Transposed Hilbert (TH).
We use PTv3 as the backbone model for all experiments, with two variants: PTv3 (Base) and PTv3 (Large). The Base variant follows the configuration used in Sonata, and the Large variant follows the configuration used in Concerto [60]. The detailed parameter settings are provided in Table 6.
@@
Table 10: Hyperparameter Analysis Experiments on the ScanNet semantic segmentation. We report Full Fine-tuning (Full-FT) and Linear Probing (LP) performance (%). Unless noted otherwise, the default configuration uses RoomTours-1k generated by $\pi^{3}$ and is pre-trained with PTv3 (Base). Both Full-FT and LP are trained for 100 epochs. Default settings are highlighted in gray.
Pre-training setting. We set the default parameters as shown in Table 7. The Laplacian smoothing loss hyperparameters were set based on experimental results and computational considerations during pre-training. In our formulation, the Laplacian smoothing loss penalizes pairwise feature differences between neighboring points on the kNN graph. When the feature difference on an edge is small, the Huber penalty behaves quadratically; when it exceeds $\delta$, it becomes linear, reducing the influence of outlier edges. Other settings related to the clustering loss follow the configuration used in Sonata. Furthermore, our comparative experiments employed three settings: RoomTours-16k (PTv3-Base), RoomTours-49k (PTv3-Base), and RoomTours-49k (PTv3-Large). Based on empirical results from previous research, the number of iterations is scaled with both data size and model size.
Downstream task setting. We evaluate the pre-training effect on two downstream tasks: semantic segmentation and instance segmentation in indoor scenes. Semantic segmentation and instance segmentation are evaluated on ScanNet [15], ScanNet200 [35], ScanNet++ [56], S3DIS [2]. For both tasks, we adopt the following fine-tuning protocols: Linear probing and Full fine-tuning. Linear probing freezes the entire backbone and trains only a linear classifier, whereas Full fine-tuning updates all model parameters. For the linear probing and the full fine-tuning protocol, we follow Sonata setting [49]. The configuration for semantic segmentation is provided in Table 8, and the configuration for instance segmentation is provided in Table 9.